OpenStudy (anonymous):

HELP again... 9+4(x+1)=25

OpenStudy (jamesj):

What do think is the first step here?

OpenStudy (jamesj):

If I were you I'd subtract 9 for each side.

OpenStudy (jamesj):

...which gives you 4(x+1) = 25 - 9 = 16 So 4(x+1) = 16 Now what?

OpenStudy (anonymous):

I'm not sure. This is tomorrow's homework, and I am not going to be in class. So I have no clue?

OpenStudy (jamesj):

Well, if I told you 4y = 12 How would you find y?

OpenStudy (anonymous):

you would divided 4/4 then 12/4 and that comes out to 3

OpenStudy (jamesj):

Right. So look now again at 4(x+1) = 16

OpenStudy (anonymous):

how did you get that? where did the 9 go?

OpenStudy (jamesj):

I started with your equation 9 + 4(x+1) = 25 and the first step was subtracting 9 from each side, as I wrote above -9 + 9 + 4(x+1) = -9 + 25 0 + 4(x+1) = 25 - 9 = 16 i.e., 4(x+1) = 16

OpenStudy (anonymous):

oh ok. I think I understand. So how did you get that into this equation? 4(x+1) = 16

OpenStudy (jamesj):

Sorry, are you asking (1) how I got to the equation 4(x+1) = 16 or (2) how now you solve this equation? The answer to number (1) are the steps I just wrote down.

OpenStudy (anonymous):

I think I know how to it. thanks :)

OpenStudy (anonymous):

now 30 more equations like this left. thanks again.

OpenStudy (jamesj):

What answer did you get?

OpenStudy (anonymous):

To 4(x+1) = 16. I got x=3

OpenStudy (jamesj):

Exactly. Great.

OpenStudy (anonymous):

now just 30 more equations. thank you :)

OpenStudy (anonymous):

one last question on this one equation 7(d-5)+12=5. So would I do 7-7(d-5)+12-7=5-7?

OpenStudy (jamesj):

No. First step: subtract 12 from both sides

OpenStudy (anonymous):

So. -5(d-5)=5

OpenStudy (jamesj):

7(d-5) + 12 = 5 7(d-5) + 12 - 12 = 5 - 12 7(d-5) + 0 = -7 That is, 7(d-5) = -7

OpenStudy (anonymous):

what now do I do? I tried to find the one we did earlier but I can not locate it.

OpenStudy (jamesj):

Now divide both sides by 7 \[\frac{7(d-5)}{7} = \frac{-7}{7}\] Cancel 7s throughout and d - 5 = -1

OpenStudy (anonymous):

d=-6? but then when I see if I am correct I come out with (-6-5)=-11?

OpenStudy (jamesj):

Noooo ..... now add 5 to both sides

OpenStudy (anonymous):

(d)=4?

OpenStudy (anonymous):

(d-0)=4?

OpenStudy (jamesj):

d - 5 = -1 d - 5 + 5 = -1 + 5 d + 0 = 4

OpenStudy (jamesj):

Yep, d = 4

OpenStudy (anonymous):

I'm so lost. I totally guessed with 4.

OpenStudy (jamesj):

Go back over these two examples and try and understand them. Copy out my steps. Then take a blank piece of paper and do it again. When you can do it without looking at this site or your previous version, then you understand them. When you understand these two problems, your other 28 problems will be much easier.